Find the difference and reduce to lowest terms ...1/3 - 1/10 .. I guessed 7/30 but I dont know how to work it out .. Can You explain how to work the problem?
Answer by Alan3354(69443)

1/3 - 1/10
-------------
Find a common denominator to add or subtract fractions.
The product always works.
3*10 = 30
============
1/3 - 1/10 = 10/30 - 3/30
= 7/30
==============
That's a good guess.